# Nonflat bands and chiral symmetry in magic-angle twisted bilayer graphene

### v1

@@ III Vortex Chern basis of the flat bands
$\begin{split}\tilde{\tau}_{z}&=\frac{i}{3\sqrt{3}}\Bigg{(}\sum_{\begin{subarray}{c}\ll\bm{r},\bm{r^{\prime}}\gg,s\\ \bm{r},\bm{r^{\prime}}\in\text{sl}\ A\end{subarray}}\eta_{\bm{rr^{\prime}}}c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}-\sum_{\begin{subarray}{c}\ll\bm{r},\bm{r^{\prime}}\gg,s\\ \bm{r},\bm{r^{\prime}}\in\text{sl}\ B\end{subarray}}\eta_{\bm{rr^{\prime}}}c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}\Bigg{)},\\ \tilde{\sigma}_{z}&=\sum_{\bm{r}\in\text{sl}\ A,s}c^{\dagger}_{\bm{r}s}c_{\bm{r}s}-\sum_{\bm{r}\in\text{sl}\ B,s}c^{\dagger}_{\bm{r}s}c_{\bm{r}s},\end{split}$ (S14)
where $\ll\bm{r},\bm{r^{\prime}}\gg$ denotes next-nearest neighbors, ’sl’ is short for sublattice and $\eta_{\bm{rr^{\prime}}}=\pm 1$ depending if the pair $\bm{r},\bm{r^{\prime}}$ is part of a clockwise or counterclockwise triangle, see Fig. S1(b). Both operators conserve the Bloch momentum and $[\tilde{\tau}_{z},\tilde{\sigma}_{z}]=0$. The valley operator is constructed so that graphene plane waves near the $K$($K^{\prime}$) point have expectation values close to $1$($-1$).
Here and throughout, single-particle operators restricted to the flat bands will be indicated by a line, $\overline{\phantom{o}}$’ (Eq. (S9) would then read $\mathcal{P}O\mathcal{P}=\overline{O}+\sum_{i}O_{ii}$). We start by simultaneously diagonalizing the flat-band matrices at each $\bm{k}$, $[\overline{\sigma}_{z}(\bm{k})]_{nm}=\langle\bm{k}n|\tilde{\sigma}_{z}|\bm{k}m\rangle$ and $[\overline{\tau}_{z}(\bm{k})]_{nm}=\langle\bm{k}n|\tilde{\tau}_{z}|\bm{k}m\rangle$. Valley $K$($K^{\prime}$) states are identified as the states with positive (negative) eigenvalue under $\overline{\tau}_{z}$. We use $\tau_{i}$, $\sigma_{i}$ for the identity ($i=0$) and Pauli ($i=x,y,z$) matrices acting in the valley and sublattice polarized sectors, respectively. Valley $K(K^{\prime})$ has eigenvalue 1(1)111(-1)1 ( - 1 ) under $\tau_{z}$, and we assign eigenvalue 1(1)111(-1)1 ( - 1 ) to sublattice $A(B)$ under $\sigma_{z}$. The resulting bands are topological with Chern numbers equal to $\sigma_{z}\tau_{z}$ [6, 4, 5].
The ’particle-hole’ operator, $\tilde{P}$, is emergent in the continuum limit of MATBG [84, 5, 54, 69]. In practice, we consider the product $C_{2z}\tilde{P}$ rather than $\tilde{P}$ (understand the product of single-particle operators in second quantization as the second-quantized version of the product of the corresponding first-quantized operators). For ease of notation we will not distinguish between the microscopic, indicated with $\tilde{\phantom{o}}$’, and projected, indicated with $\overline{\phantom{o}}$’, versions of the crystallographic and time-reversal symmetries; their distinction must be understood within the context. Note moreover that the exact symmetries are block-diagonal in the band basis. For the definition of $C_{2z}\tilde{P}$, see Section IV.
@@
$\begin{split}C_{2z}\mathcal{T}=\sigma_{x}\mathcal{K},\\ C_{2z}\overline{P}=\sigma_{y}\tau_{x},\end{split}$ (S15)
@@
As a matter of fact, the flat Hilbert space is not invariant under the action of the particle-hole operator (in other words, $C_{2z}\tilde{P}$ does not commute with the projector to the flat bands), forbidding the representation $C_{2z}\overline{P}=\sigma_{y}\tau_{x}$. This is evidenced by the singular vales of $C_{2z}\overline{P}$ in Fig. S2(a) and Table S2 (see also Refs. [54, 104]). Instead, we replace $C_{2z}\overline{P}$ by the unitary matrix $C_{2z}P=C_{2z}\overline{P}/\big{(}C_{2z}\overline{P}(C_{2z}\overline{P})^{\dagger}\big{)}^{1/2}$ and enforce Eq. (S15) for $C_{2z}P$. A similar procedure was applied when we assigned eigenvalues ±$\pm 1$ under $\sigma_{z}$ to the sublattice polarized states even though the eigenvalues of $\overline{\sigma}_{z}$ are different from ±$\pm 1$ (Fig. S2(a) and Table S2). Notice that because $C_{2z}P\neq C_{2z}\overline{P}$, $\{C_{2z}\overline{P},\overline{\sigma}_{z}\}$ need not be equal to 00 and $C_{2z}P$ might display components of the form $\tau_{x}$, $\tau_{y}$ and $\sigma_{z}\tau_{y}$ that cannot be removed with a gauge choice. Nevertheless, we find that these terms are negligible and we can satisfy Eq. (S15) up to numerical precision.
@@
$\begin{split}C_{2z}&=\begin{cases}\sigma_{x}\tau_{x}\quad&\bm{k}=\Gamma_{M}\\ -\sigma_{x}\tau_{x}\quad&\bm{k}\neq\Gamma_{M}\end{cases},\\ \mathcal{T}&=\begin{cases}\tau_{x}\mathcal{K}\quad&\bm{k}=\Gamma_{M}\\ -\tau_{x}\mathcal{K}\quad&\bm{k}\neq\Gamma_{M}\end{cases},\\ C_{3z}&=\begin{cases}\mathbb{1}\quad&\bm{k}=\Gamma_{M}\\ \text{exp}(-2\pi i/3\ \sigma_{z}\tau_{z})\quad&\bm{k}\neq\Gamma_{M}\end{cases}.\end{split}$ (S16)
@@
As is, the Chern gauge defined by Eqs. (S15),(S16) exhibits an ambiguity consisting of rotations of the states with the same Chern number, $\text{exp}(i\phi(\bm{k})\sigma_{z}\tau_{z})$. We fix this phase by requiring that the Wannier functions obtained from the Chern bands are maximally localized [87, 88], following the algorithm of Ref. [87]. By $C_{2z}\mathcal{T}$, the maximal localization procedure yields the same phase $\phi_{K}(\bm{k})$ for the $KA$ and $KB$ polarized bands, and $\phi_{K^{\prime}}(\bm{k})$ for the $K^{\prime}A$ and $K^{\prime}B$ bands. Given that there exists some particle-hole asymmetry, i.e. $C_{2z}\overline{P}\neq C_{2z}P$, $\phi_{K}(\bm{k})$ need not be equal to $\phi_{K^{\prime}}(\bm{k})$. In practice we find that they differ by at most $1^{\circ}$. We allow for this small error and take $\phi(\bm{k})$ to be the average. With this gauge transformation, the Bloch states are periodic and smooth except at the $\Gamma_{M}$ point, where the Berry connection acquires a vortex with a winding number equal to the Chern number [87, 88].
@@ IV Symmetry of MATBG
$\langle\bm{r}|\psi\rangle=\sum_{\tau}e^{i\tau\bm{K}_{l}\cdot\bm{r}}f^{\psi}_{\tau\sigma l}(\bm{r}).$ (S17)
The phases $e^{\pm i\bm{K}_{l}\cdot\bm{r}}$ oscillate rapidly on the graphene lattice, whereas the envelope functions $f_{\tau\sigma l}(\bm{r})$ vary slowly on the graphene lattice and depend on the valleys $\tau=\pm$ and on the sublattice $\sigma=A,B$ and layer $l=\text{top}\ (t),\text{bottom}\ (b)$ of the atom at $\bm{r}$. $(-)\bm{K}_{l}$ is the $K\ (K^{\prime})$ point of layer $l$, see Fig. S1. We assume that $|\psi\rangle$ has a definite spin projection that is left implicit. This decomposition is only justified for states at low energies coming from the valleys of the parent graphene monolayers. In other words, the decomposition can be written only in effective models for MATBG where remote bands above a certain cutoff have been integrated out. In continuous models, the functions $f$ defined at the atomic positions are promoted to smooth functions in the continuum.
Symmetry of ${\mathcal{H}}_{\text{int}}$.
Making use of Eq. (S17), consider the matrix element of $\mathcal{H}_{\text{int}}$ for generic states $\alpha,\beta,\gamma,\delta$ with valleys $\eta_{\alpha},\eta_{\beta},\eta_{\gamma},\eta_{\delta}$,
@@
$+U\sum_{\bm{r}}e^{i(\tau_{\delta}+\tau_{\gamma}-\tau_{\alpha}-\tau_{\beta})\bm{K}_{l}\cdot\bm{r}}f^{\alpha}_{\tau_{\alpha}\sigma l}(\bm{r})^{*}f^{\delta}_{\tau_{\delta}\sigma l}(\bm{r})f^{\beta}_{\tau_{\beta}\sigma l}(\bm{r})^{*}f^{\gamma}_{\tau_{\gamma}\sigma l}(\bm{r}).$ (S18)
The nature of the Coulomb interaction forces the implicit spin components $s_{\alpha}=s_{\delta}$, $s_{\beta}=s_{\gamma}$ for both the long range and Hubbard terms, and $s_{\alpha}=\bar{s}_{\beta}$ for the Hubbard term. Notice also that we have omitted an essentially trivial piece coming from the $-\frac{1}{2}$ subtraction, see Eq. (S2).
Clearly, the matrix element of Eq. (S18) is the same as ${\mathcal{H}}_{\text{int},\alpha\beta\gamma\delta}$ if the initial and final states belong to the central bands. Nonetheless, ${\mathcal{H}}_{\text{int}}$ includes also a nontrivial local potential
$\mathcal{H}_{\text{int}}\supset\sum_{\bm{r}s}\pi(\bm{r})\rho_{\bm{r}s},\quad\quad\quad\quad\quad\pi(\bm{r})=-2\sum_{\bm{r^{\prime}}\neq\bm{r}}V(\bm{r}-\bm{r^{\prime}})\rho_{0\bm{r^{\prime}}}-U\rho_{0\bm{r}},$ (S19)
see Eqs. (S12), (S13), that we ought to consider separately.
Valley-spin flavor symmetries.
Looking at the first line of Eq. (S18), only if $\tau_{\alpha}=\tau_{\delta}$ and $\tau_{\beta}=\tau_{\gamma}$ the matrix elements is non vanishing. Indeed, if $\tau_{\delta}=-\tau_{\alpha}$, we can take the $\bm{r}$, $\bm{r}+\bm{a}_{1}$ and $\bm{r}+\bm{a}_{2}$ and compute the phases $e^{2i\tau_{\delta}\bm{K}_{l}\cdot\bm{r}}$, $e^{2i\tau_{\delta}\bm{K}_{l}\cdot(\bm{r}+\bm{a}_{1}/2)}=e^{2i\tau_{\delta}\bm{K}_{l}\cdot\bm{r}}e^{\pm 2\pi i/3}$ and $e^{2i\tau_{\delta}\bm{K}_{l}\cdot(\bm{r}+\bm{a}_{2}/2)}=e^{2i\tau_{\delta}\bm{K}_{l}\cdot\bm{r}}e^{\mp 2\pi i/3}$. Given that the $f$ functions and the potential $V(\bm{r})$ vary slowly on atomic distances, $V(\bm{r}+\bm{a}_{1,2})\approx V(\bm{r})$, $f(\bm{r}+\bm{a}_{1,2})\approx f(\bm{r})$, the sum over the corresponding three atoms in ${\mathcal{H}}_{\text{int},\alpha\beta\gamma\delta}$ (modulo lattice relaxation and twist) cancels by $01+e^{\pm 2\pi i/3}+e^{\mp 2\pi i/3}=$ 0. Similarly occurs if $\tau_{\gamma}=-\tau_{\beta}$.
@@
$\big{[}s_{i},{\mathcal{H}}_{\text{int}}\big{]}=\big{[}s_{i}\tau_{z},{\mathcal{H}}_{\text{int}}\big{]}=0,\quad\quad\quad\quad\quad(i=0,x,y,z).$ (S20)
@@
With respect to the second line of Eq. (S18), by a similar argument we deduce that valley charge is conserved, $\tau_{\alpha}+\tau_{\beta}=\tau_{\gamma}+\tau_{\delta}$. However, only global spin rotations preserve the Hubbard Hamiltonian, breaking $U(2)_{K}\times U(2)_{K^{\prime}}$ down to the physical symmetry of $SU(2)_{\text{spin}}\times U(1)_{\text{V}}$. Furthermore, the long range interaction also breaks $U(2)_{K}\times U(2)_{K^{\prime}}$. Indeed, in the argument presented previously we assumed that the the potential $V(\bm{r})$ is slowly varying on atomic distances. This is not true near $\bm{r}=0$ where there is a steep $1/|\bm{r}|$ dependence. In turn, there are symmetry breaking components coming from small distances. $U(2)_{K}\times U(2)_{K^{\prime}}$-breaking interactions are called ’intervalley Hund’s’ couplings.
@@
Now, consider the ’chiral’ operator $\tilde{\sigma}_{z}$ in Eq. (S14). The exponential $\text{exp}(i\varphi\tilde{\sigma}_{z})$ attaches the phase $e^{i\varphi}$ to $f_{\tau Al}$ and $e^{-i\varphi}$ to $f_{\tau Bl}$. Clearly, this is an exact symmetry of ${{H}}_{\text{int}}$ because each phase encounters the complex conjugate in the matrix element, Eq. (S18).
Nonetheless, the mode given by $\text{exp}(i\varphi\tilde{\sigma}_{z})|\alpha\rangle$ generically does not belong to the subspace of central bands (in other words, it is not possible in general to construct a perfectly sublattice polarized basis in the effective theory), spoiling the chiral symmetry in $\mathcal{H}_{\text{int}}$
To overcome this problem, we redefine the sublattice operator on the central bands as $\sigma_{z}=\sum_{\alpha\beta}[\sigma_{z}]_{\alpha\beta}c^{\dagger}_{\alpha}c_{\beta}$, with the matrix $\sigma_{z}=\overline{\sigma}_{z}/\big{(}\overline{\sigma}_{z}\overline{\sigma}_{z}^{\dagger}\big{)}^{1/2}$ and $[\overline{\sigma}_{z}]_{\alpha\beta}=\langle\alpha|\tilde{\sigma}_{z}|\beta\rangle$. This $\sigma_{z}$ here is the same as $\sigma_{z}$ acting on the Chern basis defined below Eq. (S14), having the property that it is the unitary matrix closest to $\overline{\sigma}_{z}$ (in the norm $||M||=\text{tr}(M^{\dagger}M)$). We can assess the discrepancy between $\overline{\sigma}_{z}$ and $\sigma_{z}$ via the singular values or, equivalently up to a sign, the eigenvalues of $\overline{\sigma}_{z}$, see Table (S2). Singular values equal to $1$ imply that $\sigma_{z}=\overline{\sigma}_{z}$.
With respect to the local potential, in the perfectly sublattice polarized limit the symmetry is exact owing to $\sigma_{z}\rho_{\bm{r}z}\sigma_{z}=\rho_{\bm{r}s}$. However, from the fact that $\overline{\sigma}_{z}\neq\sigma_{z}$ there appear intersublattice scatterings introducing some amount of symmetry breaking.
@@
$\big{[}\sigma_{z}\tau_{z},{\mathcal{H}}_{\text{int}}\big{]}\approx 0,$ (S21)
@@
Particle-hole symmetry.
Let us turn to the particle-hole operator, $C_{2z}\tilde{P}$. This is a single-particle operator, local in real space and defined by its action on the $f$ functions in Eq. (S17),
$C_{2z}\tilde{P}=\sigma_{x}\tau_{y}\mu_{y},$ (S22)

### v3

@@ Appendix C Vortex Chern basis of the flat bands
$\begin{split}\tilde{\tau}_{z}&=\frac{i}{3\sqrt{3}}\Bigg{(}\sum_{\begin{subarray}{c}\ll\bm{r},\bm{r^{\prime}}\gg,s\\ \bm{r},\bm{r^{\prime}}\in\text{sl}\ A\end{subarray}}\eta_{\bm{rr^{\prime}}}c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}-\sum_{\begin{subarray}{c}\ll\bm{r},\bm{r^{\prime}}\gg,s\\ \bm{r},\bm{r^{\prime}}\in\text{sl}\ B\end{subarray}}\eta_{\bm{rr^{\prime}}}c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}\Bigg{)},\\ \tilde{\sigma}_{z}&=\sum_{\bm{r}\in\text{sl}\ A,s}c^{\dagger}_{\bm{r}s}c_{\bm{r}s}-\sum_{\bm{r}\in\text{sl}\ B,s}c^{\dagger}_{\bm{r}s}c_{\bm{r}s},\end{split}$ (18)
where $\ll\bm{r},\bm{r^{\prime}}\gg$ denotes next-nearest neighbors, ’sl’ is short for sublattice and $\eta_{\bm{rr^{\prime}}}=\pm 1$ depending if the pair $\bm{r},\bm{r^{\prime}}$ is part of a clockwise or counterclockwise triangle, see Fig. A1(b). Both operators conserve the Bloch momentum and $[\tilde{\tau}_{z},\tilde{\sigma}_{z}]=0$. The valley operator is constructed so that graphene plane waves near the $K$($K^{\prime}$) point have expectation values close to $1$($-1$).
Here and throughout, single-particle operators restricted to the flat bands will be indicated by a line, $\overline{\phantom{o}}$’ (Eq. (13) would then read $\mathcal{P}O\mathcal{P}=\overline{O}+\sum_{i}O_{ii}$). We start by simultaneously diagonalizing the flat-band matrices at each $\bm{k}$, $[\overline{\sigma}_{z}(\bm{k})]_{nm}=\langle\bm{k}n|\tilde{\sigma}_{z}|\bm{k}m\rangle$ and $[\overline{\tau}_{z}(\bm{k})]_{nm}=\langle\bm{k}n|\tilde{\tau}_{z}|\bm{k}m\rangle$. Valley $K$($K^{\prime}$) states are identified as the states with positive (negative) eigenvalue under $\overline{\tau}_{z}$. We use $\tau_{i}$, $\sigma_{i}$ for the identity ($i=0$) and Pauli ($i=x,y,z$) matrices acting in the valley and sublattice polarized sectors, respectively. Valley $K(K^{\prime})$ has eigenvalue 1(1)111(-1)1 ( - 1 ) under $\tau_{z}$, and we assign eigenvalue 1(1)111(-1)1 ( - 1 ) to sublattice $A(B)$ under $\sigma_{z}$. The resulting bands are topological with Chern numbers equal to $\sigma_{z}\tau_{z}$ [6, 4, 5].
The ’particle-hole’ operator, $\tilde{P}$, is emergent in the continuum limit of MATBG [85, 5, 56, 71]. In practice, we consider the product $C_{2z}\tilde{P}$ rather than $\tilde{P}$ (understand the product of single-particle operators in second quantization as the second-quantized version of the product of the corresponding first-quantized operators). For ease of notation we will not distinguish between the microscopic, indicated with $\tilde{\phantom{o}}$’, and projected, indicated with $\overline{\phantom{o}}$’, versions of the crystallographic and time-reversal symmetries; their distinction must be understood within the context. Note moreover that the exact symmetries are block-diagonal in the band basis. For the definition of $C_{2z}\tilde{P}$, see Appendix D.
@@
$\begin{split}C_{2z}\mathcal{T}=\sigma_{x}\mathcal{K},\\ C_{2z}\overline{P}=\sigma_{y}\tau_{x},\end{split}$ (19)
@@
As a matter of fact, the flat Hilbert space is not invariant under the action of the particle-hole operator (in other words, $C_{2z}\tilde{P}$ does not commute with the projector to the flat bands), forbidding the representation $C_{2z}\overline{P}=\sigma_{y}\tau_{x}$. This is evidenced by the singular vales of $C_{2z}\overline{P}$ in Fig. D2(a) and Table D2 (see also Refs. [56, 99]). Instead, we replace $C_{2z}\overline{P}$ by the unitary matrix $C_{2z}P=C_{2z}\overline{P}/\big{(}C_{2z}\overline{P}(C_{2z}\overline{P})^{\dagger}\big{)}^{1/2}$ and enforce Eq. (19) for $C_{2z}P$. A similar procedure was applied when we assigned eigenvalues ±$\pm 1$ under $\sigma_{z}$ to the sublattice polarized states even though the eigenvalues of $\overline{\sigma}_{z}$ are different from ±$\pm 1$ (Fig. D2(a) and Table D2). Notice that because $C_{2z}P\neq C_{2z}\overline{P}$, $\{C_{2z}\overline{P},\overline{\sigma}_{z}\}$ need not be equal to 00 and $C_{2z}P$ might display components of the form $\tau_{x}$, $\tau_{y}$ and $\sigma_{z}\tau_{y}$ that cannot be removed with a gauge choice. Nevertheless, we find that these terms are negligible and we can satisfy Eq. (19) up to numerical precision.
@@
$\begin{split}C_{2z}&=\begin{cases}\sigma_{x}\tau_{x}\quad&\bm{k}=\Gamma_{M}\\ -\sigma_{x}\tau_{x}\quad&\bm{k}\neq\Gamma_{M}\end{cases},\\ \mathcal{T}&=\begin{cases}\tau_{x}\mathcal{K}\quad&\bm{k}=\Gamma_{M}\\ -\tau_{x}\mathcal{K}\quad&\bm{k}\neq\Gamma_{M}\end{cases},\\ C_{3z}&=\begin{cases}\mathbb{1}\quad&\bm{k}=\Gamma_{M}\\ \text{exp}(-2\pi i/3\ \sigma_{z}\tau_{z})\quad&\bm{k}\neq\Gamma_{M}\end{cases}.\end{split}$ (20)
@@
As is, the Chern gauge defined by Eqs. (19),(20) exhibits an ambiguity consisting of rotations of the states with the same Chern number, $\text{exp}(i\phi(\bm{k})\sigma_{z}\tau_{z})$. We fix this phase by requiring that the Wannier functions obtained from the Chern bands are maximally localized [88, 89], following the algorithm of Ref. [88]. By $C_{2z}\mathcal{T}$, the maximal localization procedure yields the same phase $\phi_{K}(\bm{k})$ for the $KA$ and $KB$ polarized bands, and $\phi_{K^{\prime}}(\bm{k})$ for the $K^{\prime}A$ and $K^{\prime}B$ bands. Given that there exists some particle-hole asymmetry, i.e. $C_{2z}\overline{P}\neq C_{2z}P$, $\phi_{K}(\bm{k})$ need not be equal to $\phi_{K^{\prime}}(\bm{k})$. In practice we find that they differ by at most $1^{\circ}$. We allow for this small error and take $\phi(\bm{k})$ to be the average. With this gauge transformation, the Bloch states are periodic and smooth except at the $\Gamma_{M}$ point, where the Berry connection acquires a vortex with a winding number equal to the Chern number [88, 89].
@@ Appendix D Symmetry of MATBG
$\langle\bm{r}|\psi\rangle=\sum_{\tau}e^{i\tau\bm{K}_{l}\cdot\bm{r}}f^{\psi}_{\tau\sigma l}(\bm{r}).$ (21)
The phases $e^{\pm i\bm{K}_{l}\cdot\bm{r}}$ oscillate rapidly on the graphene lattice, whereas the envelope functions $f_{\tau\sigma l}(\bm{r})$ vary slowly on the graphene lattice and depend on the valleys $\tau=\pm$ and on the sublattice $\sigma=A,B$ and layer $l=\text{top}\ (t),\text{bottom}\ (b)$ of the atom at $\bm{r}$. $(-)\bm{K}_{l}$ is the $K\ (K^{\prime})$ point of layer $l$, see Fig. A1. We assume that $|\psi\rangle$ has a definite spin projection that is left implicit. This decomposition is only justified for states at low energies coming from the valleys of the parent graphene monolayers. In other words, the decomposition can be written only in effective models for MATBG where remote bands above a certain cutoff have been integrated out. In continuous models, the functions $f$ defined at the atomic positions are promoted to smooth functions in the continuum.
Symmetry of ${\mathcal{H}}_{\text{int}}$
Making use of Eq. (21), consider the matrix element of $\mathcal{H}_{\text{int}}$ for generic states $\alpha,\beta,\gamma,\delta$ with valleys $\eta_{\alpha},\eta_{\beta},\eta_{\gamma},\eta_{\delta}$,
@@
$+U\sum_{\bm{r}}e^{i(\tau_{\delta}+\tau_{\gamma}-\tau_{\alpha}-\tau_{\beta})\bm{K}_{l}\cdot\bm{r}}f^{\alpha}_{\tau_{\alpha}\sigma l}(\bm{r})^{*}f^{\delta}_{\tau_{\delta}\sigma l}(\bm{r})f^{\beta}_{\tau_{\beta}\sigma l}(\bm{r})^{*}f^{\gamma}_{\tau_{\gamma}\sigma l}(\bm{r}).$ (22)
The nature of the Coulomb interaction forces the implicit spin components $s_{\alpha}=s_{\delta}$, $s_{\beta}=s_{\gamma}$ for both the long range and Hubbard terms, and $s_{\alpha}=\bar{s}_{\beta}$ for the Hubbard term. Notice also that we have omitted an essentially trivial piece coming from the $-\frac{1}{2}$ subtraction, see Eq. (6).
Clearly, the matrix element of Eq. (22) is the same as ${\mathcal{H}}_{\text{int},\alpha\beta\gamma\delta}$ if the initial and final states belong to the central bands. Nonetheless, ${\mathcal{H}}_{\text{int}}$ includes also a nontrivial local potential
$\mathcal{H}_{\text{int}}\supset\sum_{\bm{r}s}\pi(\bm{r})\rho_{\bm{r}s},\quad\quad\quad\quad\quad\pi(\bm{r})=-2\sum_{\bm{r^{\prime}}\neq\bm{r}}V(\bm{r}-\bm{r^{\prime}})\rho_{0\bm{r^{\prime}}}-U\rho_{0\bm{r}},$ (23)
see Eqs. (16), (17), that we ought to consider separately.
Valley-spin symmetries.
Looking at the first line of Eq. (22), only if $\tau_{\alpha}=\tau_{\delta}$ and $\tau_{\beta}=\tau_{\gamma}$ the matrix elements is non vanishing. Indeed, if $\tau_{\delta}=-\tau_{\alpha}$, we can take the $\bm{r}$, $\bm{r}+\bm{a}_{1}$ and $\bm{r}+\bm{a}_{2}$ and compute the phases $e^{2i\tau_{\delta}\bm{K}_{l}\cdot\bm{r}}$, $e^{2i\tau_{\delta}\bm{K}_{l}\cdot(\bm{r}+\bm{a}_{1}/2)}=e^{2i\tau_{\delta}\bm{K}_{l}\cdot\bm{r}}e^{\pm 2\pi i/3}$ and $e^{2i\tau_{\delta}\bm{K}_{l}\cdot(\bm{r}+\bm{a}_{2}/2)}=e^{2i\tau_{\delta}\bm{K}_{l}\cdot\bm{r}}e^{\mp 2\pi i/3}$. Given that the $f$ functions and the potential $V(\bm{r})$ vary slowly on atomic distances, $V(\bm{r}+\bm{a}_{1,2})\approx V(\bm{r})$, $f(\bm{r}+\bm{a}_{1,2})\approx f(\bm{r})$, the sum over the corresponding three atoms in ${\mathcal{H}}_{\text{int},\alpha\beta\gamma\delta}$ (modulo lattice relaxation and twist) cancels by $01+e^{\pm 2\pi i/3}+e^{\mp 2\pi i/3}=$ 0. Similarly occurs if $\tau_{\gamma}=-\tau_{\beta}$.
@@
$\big{[}s_{i},{\mathcal{H}}_{\text{int}}\big{]}=\big{[}s_{i}\tau_{z},{\mathcal{H}}_{\text{int}}\big{]}=0,\quad\quad\quad\quad\quad(i=0,x,y,z).$ (24)
@@
With respect to the second line of Eq. (22), by a similar argument we deduce that valley charge is conserved, $\tau_{\alpha}+\tau_{\beta}=\tau_{\gamma}+\tau_{\delta}$. However, only global spin rotations preserve the Hubbard Hamiltonian, breaking $U(2)_{K}\times U(2)_{K^{\prime}}$ down to the physical symmetry of $SU(2)_{\text{spin}}\times U(1)_{\text{V}}$. Furthermore, the long range interaction also breaks $U(2)_{K}\times U(2)_{K^{\prime}}$. Indeed, in the argument presented previously we assumed that the the potential $V(\bm{r})$ is slowly varying on atomic distances. This is not true near $\bm{r}=0$ where there is a steep $1/|\bm{r}|$ dependence. In turn, there are symmetry breaking components coming from small distances. $U(2)_{K}\times U(2)_{K^{\prime}}$-breaking interactions are called ’intervalley Hund’s’ couplings.
@@
Now, consider the ’chiral’ operator $\tilde{\sigma}_{z}$ in Eq. (18). The exponential $\text{exp}(i\varphi\tilde{\sigma}_{z})$ attaches the phase $e^{i\varphi}$ to $f_{\tau Al}$ and $e^{-i\varphi}$ to $f_{\tau Bl}$. Clearly, this is an exact symmetry of ${{H}}_{\text{int}}$ because each phase encounters the complex conjugate in the matrix element, Eq. (22).
Nonetheless, the mode given by $\text{exp}(i\varphi\tilde{\sigma}_{z})|\alpha\rangle$ generically does not belong to the subspace of central bands (in other words, it is not possible in general to construct a perfectly sublattice polarized basis in the effective theory), spoiling the chiral symmetry in $\mathcal{H}_{\text{int}}$.
To overcome this problem, we redefine the sublattice operator on the central bands as $\sigma_{z}=\sum_{\alpha\beta}[\sigma_{z}]_{\alpha\beta}c^{\dagger}_{\alpha}c_{\beta}$, with the matrix $\sigma_{z}=\overline{\sigma}_{z}/\big{(}\overline{\sigma}_{z}\overline{\sigma}_{z}^{\dagger}\big{)}^{1/2}$ and $[\overline{\sigma}_{z}]_{\alpha\beta}=\langle\alpha|\tilde{\sigma}_{z}|\beta\rangle$. This $\sigma_{z}$ here is the same as $\sigma_{z}$ acting on the Chern basis defined below Eq. (18), having the property that it is the unitary matrix closest to $\overline{\sigma}_{z}$ (in the norm $||M||=\text{tr}(M^{\dagger}M)$). We can assess the discrepancy between $\overline{\sigma}_{z}$ and $\sigma_{z}$ via the singular values or, equivalently up to a sign, the eigenvalues of $\overline{\sigma}_{z}$, see Table (D2). Singular values equal to $1$ imply that $\sigma_{z}=\overline{\sigma}_{z}$.
With respect to the local potential, in the perfectly sublattice polarized limit the symmetry is exact owing to $\sigma_{z}\rho_{\bm{r}s}\sigma_{z}=\rho_{\bm{r}s}$. However, from the fact that $\overline{\sigma}_{z}\neq\sigma_{z}$ there appear intersublattice scatterings introducing some amount of symmetry breaking.
@@
$\big{[}\sigma_{z}\tau_{z},{\mathcal{H}}_{\text{int}}\big{]}\approx 0,$ (25)
@@
$C_{2z}P$ symmetry.
Let us turn to the particle-hole operator, $C_{2z}\tilde{P}$. This is a single-particle operator, local in real space and defined by its action on the $f$ functions in Eq. (21),
$C_{2z}\tilde{P}=\sigma_{x}\tau_{y}\mu_{y},$ (26)